# Detecting Near-Duplicate Face Images

### v1

@@ 3.3 Link Prediction
The link prediction module accepts a vector of depth labels $\bm{l}\in\mathbb{R}^{M}$ for a set of $M$ near-duplicate images and outputs a data structure containing a set of vertices $V$, such that $|V|=M$, and directed edges $E$. We refer this directed structure as the image phylogeny tree: $\bm{IPT}(V,E)$. Let $g(\cdot,\cdot)$ represent the link prediction module that requires a pair of inputs $\bm{l}$ and $\bm{N}$. Therefore, $\bm{IPT}=g(\bm{l},\bm{N})$. Here, $\bm{l}$ refers to the depth labels computed using the node embedding step. $\bm{N}$ represents the sensor noise pattern features computed from the images $\bm{I}$, such that, $\bm{N}\in\mathbb{R}^{M\times d^{2}}$. The link prediction performs two tasks: (i) depth label correction, and (ii) pairwise link inference.
@@
2Extract the sensor noise pattern (PRNU) for each image $\bm{N}(i)$ where $i=1,\cdots,M$
@@
For depth label correction, we first check whether multiple nodes have depth label=1, because an IPT can have only one root node. Multiple root nodes exist if $|Candidateroots|>1$, where, $Candidateroots=arg_{i}\{\bm{l}(i)==1\}$, where $i=1,\cdots,M$. We employed sensor noise pattern features present in images for depth label correction. Photo Response Non-Uniformity (PRNU) is a type of sensor pattern noise which manifests in the image as a result of non-uniform response of the pixels to the same light intensity [41], and is typically used for sensor identification. Photometric and geometric transformations can induce changes in the PRNU pattern present in an image [41, 42]. See Fig. 6 to visualize how PRNU patterns change in presence of transformations. We used the power spectral density (PSD) plots to demonstrate variations in PRNU patterns that can help discriminate between original and transformed images. Firstly, we compute the Euclidean distance between the PRNU features of each of the candidate root nodes and the remaining nodes, $\bm{D}_{c}=\|(\bm{N}(Candidateroots(c))-\bm{N}(z))\|^{2}_{2}$, where $1\leq z\leq M,z\neq c$, $\bm{N}$ corresponds to PRNU for each image. As the depth increases, it implies that the root node (original image) has undergone multiple sequences of transformations, resulting in higher variation in PRNU patterns. So, we retain that node which results in the highest distance as the correct root node, $l(c)=1$, where, $c=argmax(\bm{D})$. Finally, we deduce the root node as the node with corrected depth label = 1, $R=\arg(\bm{l}_{corr}(i)==1),1\leq i\leq M$, where, $\bm{l}_{corr}$ represents the set of corrected depth labels. The remaining nodes that were misclassified as depth label=1 are then re-assigned to depth label=2. The reason for re-assignment to depth label=2 is simply because higher depth labels would imply missing intermediate nodes. So, to avoid ambiguity we assign them to depth label=2.
@@
Figure 6: PRNU in image phylogeny. The variations in PRNU are better visualized using the binary maps (threshold=0) and their corresponding power spectral density (PSD) plots. Note that the PSD plots of the PRNU patterns reveal discernible differences compared to the images.
@@ 5 Datasets and Experiments
In this experiment, we evaluated the proposed method in terms of (i) root identification accuracy, which computes the proportion of correctly identified root nodes, and (ii) IPT reconstruction accuracy, which is computed as follows: $\frac{|\textit{Original\textunderscore edges}\cap\textit{Reconstructed\textunderscore edges}|}{|\textit{Original\textunderscore edges}|}$ for face images subjected to four photometric (Brightness adjustment, Gamma transformation, Median filtering and Gaussian smoothing) and three geometric (Rotation, Scaling and Translation) transformations. We used face images from the Labeled Faces in the Wild (LFW) dataset [46]. All the images are cropped to a fixed size of 96 $\times$ 96 using a commercial face SDK. See Table I for the parameters used in generating the near-duplicates. For the node embedding module, we used a training set of 1,500 IPTs generated from 7,500 face images belonging to 123 subjects, and a validation set of 600 IPTs generated from 3,000 face images belonging to 46 subjects. There can be $n^{(n-1)}$ non-isomorphic IPTs with $n$ nodes (using Cayley’s formula [47] $n\times n^{(n-2)}$). We selected six configurations as depicted in Fig. 7 for training the graph neural networks, as these configurations cover maximum breadth and depth values possible for an IPT with five nodes.
@@ 6 Findings and Analysis
In the context of demographic variations, we reported the results for five different demographic groups (Asian, Black, Hispanic, Indian and White) in terms of root identification accuracy at Rank 2 and IPT reconstruction accuracy. The reason for using Rank 2 root identification accuracy is that we observed that Rank 2 root identification accuracy is $\sim$100for both Experiment 1 and Experiment 2, i.e., we are almost guaranteed to find root node at Rank 2. The top-$k$ root nodes can be identified as described in Sec. 3.3. We achieved 100% root identification accuracy at Rank 2 across all five demographic groups. We achieved an IPT reconstruction accuracy of 71.4% for Asian subjects, 68.6% for Black subjects, 66.1% for Hispanic subjects, 71.2% for Indian subjects and 71.8% for White subjects. The performance indicates strong robustness of the proposed method across different ethnic groups, even though the graph neural network is trained primarily on White subjects. This demonstrates that the method learns the global and local relationship irrespective of the scene content.
@@ 6.5 Ablation
TABLE VI: Comparison between pixel intensity, sensor noise pattern and convolutional features: Inception-ResNet-V2 (IRNet) and ResNet100-ArcFace (ArcFace) on the root identification accuracy on the CelebAHQ-FM dataset.
@@
Impact of convolutional feature maps as inputs to graph neural network. The graph neural network, $f(\mathbf{X},\mathbf{A})$ accepts a pair of inputs: features of the nodes ($\mathbf{X}$) and the adjacency matrix ($\mathbf{A}$). The adjacency matrix construction also requires features. We ablated using convolutional feature maps in conjunction with pixel intensity feature values and sensor pattern noise features. Therefore, we examined three different scenarios: (1) $\mathbf{X}\leftarrow$ pixel values, $\mathbf{A}\leftarrow$ sensor pattern noise, (2) $\mathbf{X}\leftarrow$ pixel values, $\mathbf{A}\leftarrow$ convolutional features and (3) $\mathbf{X}\leftarrow$ convolutional features, $\mathbf{A}\leftarrow$ sensor pattern noise. Pixel intensity values are simply flattened intensity values of the face images (we only used the first channel and resized them to 96$\times$96 for compactness); sensor noise pattern values are computed using the Enhanced PRNU features [39], and convolutional features are computed using two networks: (a) Inception-ResNet-V2 [58] that has 824 layers, and trained on the ImageNet dataset. We use the features from the final average pooling layer corresponding to $\mathbb{R}^{1\times 1536}$ embedding for each image. (b) ResNet100-ArcFace [59] as the face recognition model trained with ArcFace loss on the Glint360K face dataset from the InsightFace package with antelopev2 model. We used the output features of the model corresponding to $\mathbb{R}^{1\times 512}$ embedding for each image. Refer to the results obtained in Table VI.
@@
Comparison with image provenance algorithms. For this ablation study, we compared our performance with [19]. Although their objective was different than ours, their graph construction for the near-duplicates can be used as baseline against our method. Therefore, we adapted the official code provided by the original authors  [^8] https://github.com/CVRL/Scalable_Provenance/tree/master and compared with the graph building framework used in their paper. The baseline method performs the following steps. (i) GCM-based dissimilarity calculation in which they perform interest point detection on a set of near-duplicates followed by geometrically consistent matching (GCM) of the interest points. (ii) MI-based dissimilarity calculation in which they perform homography estimation from the GCM followed by mutual information (MI) calculation from the warped and color-matched patches. (iii) Construction of provenance graph in which they perform Kruskal’s Minimal Spanning Tree algorithm using the dissimilarity matrix. Refer to [19] for details. We used the SURF descriptor and similarity transformation for homography estimation. As the dissimilarity matrix from MI was asymmetric, we used the Oriented Kruskal [10] algorithm (classical Kruskal’s algorithm needs symmetric matrix). We achieved root identification and IPT reconstruction accuracy $\sim 0.1\%$. We tried different combinations of using only GCM-dissimilarity matrix, MI-dissimilarity matrix and combined (min-max normalized) dissimilarity matrix. However, we did not observe improvement in performance. We analyzed the matrices and present examples of MI-based dissimilarity matrix on a set of 13 near-duplicate test set (before normalization) below. Note MI captures similarity between two images, the authors in [19] use it to determine relationship between the query and the donor images that were used to generate the composite. However, in our scenario with near-duplicate faces with exactly same pose and background, implying source and target images are perfectly registered, mutual information is not discriminative enough to deduce the relationship between the near-duplicates. The authors indicated if $d_{ij}>d_{ji}$, then $I_{i}$ probably generated $I_{j}$. But we observe in most cases, the row corresponding to Ori (original image) contains low values of MI compared to column which is incorrect as we know that the edited images are derived from the original. Our findings reaffirm that mutual information and geometric consistent matching may work on composite images originating from different donor images but fails to distinguish between subtle edits in face images.

### v2

@@ 3.3 Link Prediction
The link prediction module accepts a vector of depth labels $\bm{l}\in\mathbb{R}^{M}$ for a set of $M$ near-duplicate images and outputs a data structure containing a set of vertices $V$, such that $|V|=M$, and directed edges $E$. We refer this directed structure as the image phylogeny tree: $\bm{IPT}(V,E)$. Let $g(\cdot,\cdot)$ represent the link prediction module that requires a pair of inputs $\bm{l}$ and $\bm{N}$. Therefore, $\bm{IPT}=g(\bm{l},\bm{N})$. Here, $\bm{l}$ refers to the depth labels computed using the node embedding step. $\bm{N}$ represents the sensor pattern noise features computed from the images $\bm{I}$, such that, $\bm{N}\in\mathbb{R}^{M\times d^{2}}$. The link prediction performs two tasks: (i) depth label correction, and (ii) pairwise link inference.
@@
2Extract the sensor pattern noise (PRNU) for each image $\bm{N}(i)$ where $i=1,\cdots,M$
@@
For depth label correction, we first check whether multiple nodes have depth label=1, because an IPT can have only one root node. Multiple root nodes exist if $|Candidateroots|>1$, where, $Candidateroots=arg_{i}\{\bm{l}(i)==1\}$, where $i=1,\cdots,M$. We employed sensor pattern noise features present in images for depth label correction. Photo Response Non-Uniformity (PRNU) is a type of sensor pattern noise which manifests in the image as a result of non-uniform response of the pixels to the same light intensity [42], and is typically used for sensor identification. Photometric and geometric transformations can induce changes in the PRNU pattern present in an image [42, 43]. See Fig. 6 to visualize how PRNU patterns change in presence of transformations. We used the power spectral density (PSD) plots to demonstrate variations in PRNU patterns that can help discriminate between original and transformed images. Firstly, we compute the Euclidean distance between the PRNU features of each of the candidate root nodes and the remaining nodes, $\bm{D}_{c}=\|(\bm{N}(Candidateroots(c))-\bm{N}(z))\|^{2}_{2}$, where $1\leq z\leq M,z\neq c$, $\bm{N}$ corresponds to PRNU for each image. As the depth increases, it implies that the root node (original image) has undergone multiple sequences of transformations, resulting in higher variation in PRNU patterns. So, we retain that node which results in the highest distance as the correct root node, $l(c)=1$, where, $c=argmax(\bm{D})$. Finally, we deduce the root node as the node with corrected depth label = 1, $R=\arg(\bm{l}_{corr}(i)==1),1\leq i\leq M$, where, $\bm{l}_{corr}$ represents the set of corrected depth labels. The remaining nodes that were misclassified as depth label=1 are then re-assigned to depth label=2. The reason for re-assignment to depth label=2 is simply because higher depth labels would imply missing intermediate nodes. So, to avoid ambiguity we assign them to depth label=2.
@@
Feature Selection. Our method performs the first step of clustering or grouping to correctly identify the number of trees in the forest. Images captured by different sensors at different times should be assigned to different groups. After grouping, the node embedding step performs depth label prediction for the near-duplicates in each group. The number of features needed depends on the complexity of the task. Both spectral clustering and node embedding steps analyze all images globally that require multiple rich features such as face descriptors, pixel intensity values and sensor pattern noise. Finally, link prediction infers edge/link direction between a pair of nodes locally. Convolutional features can be agnostic to photometric transformations, while pixel intensity is not sufficient to distinguish between parent (original) and child (transformed image). Hence, we use PRNU as it can effectively distinguish between parent and child nodes.
Figure 6: Photo Response Non-Uniformity (PRNU) signatures for three near-duplicates are presented. The variations in the PRNU patterns (grayscale images) are better visualized using the binary maps (threshold=0) and their corresponding power spectral density (PSD) plots. Note that the PSD plots of the PRNU patterns reveal discernible differences compared to the images demonstrating that PRNU effectively distinguishes between near-duplicates.
@@ 5 Datasets and Experiments
In this experiment, we evaluated the proposed method in terms of (i) root identification accuracy, which computes the proportion of correctly identified root nodes, and (ii) IPT reconstruction accuracy, which is computed as follows: $\frac{|\textit{Original\textunderscore edges}\cap\textit{Reconstructed\textunderscore edges}|}{|\textit{Original\textunderscore edges}|}$ for face images subjected to four photometric (Brightness adjustment, Gamma transformation, Median filtering and Gaussian smoothing) and three geometric (Rotation, Scaling and Translation) transformations. We used face images from the Labeled Faces in the Wild (LFW) dataset [47]. All the images are cropped to a fixed size of 96 $\times$ 96 using a commercial face SDK. See Table I for the parameters used in generating the near-duplicates. For the node embedding module, we used a training set of 1,500 IPTs generated from 7,500 face images belonging to 123 subjects, and a validation set of 600 IPTs generated from 3,000 face images belonging to 46 subjects. The number of labeled spanning trees with $n$ nodes is $n^{(n-2)}$ using Cayley’s formula [48]. We selected six configurations as depicted in Fig. 7 for training the graph neural networks, as these configurations cover maximum breadth and depth values possible for an IPT with five nodes.
@@ 6 Findings and Analysis
In the context of demographic variations, we reported the results for five different demographic groups (Asian, Black, Hispanic, Indian and White) in terms of root identification accuracy at Rank 2 and IPT reconstruction accuracy. The reason for using Rank 2 is that we observed that Rank 2 accuracy is $\sim$100for both Experiment 1 and Experiment 2, i.e., we are guaranteed to find the correct root node at Rank 2. Therefore, we want to focus on the performance variation across demographic groups by benchmarking it using the best use-case of our method. The top-$k$ root nodes can be identified as described in Sec. 3.3. We achieved 100% root identification accuracy at Rank 2 across all five demographic groups. We achieved an IPT reconstruction accuracy of 71.4% for Asian subjects, 68.6% for Black subjects, 66.1% for Hispanic subjects, 71.2% for Indian subjects and 71.8% for White subjects. The performance indicates strong robustness of the proposed method across different ethnic groups, even though the graph neural network is trained primarily on White subjects. This demonstrates that the method learns the global and local relationship irrespective of the scene content.
@@ 6.5 Ablation
TABLE VI: Comparison between pixel, PRNU and convolutional features on the root identification accuracy on the CelebAHQ-FM dataset. The graph neural network requires features of the nodes ($\mathbf{X}$) and the adjacency matrix ($\mathbf{A}$). We utilize two types of convolutional features in our work, namely, Inception-ResNet-V2 (IRNet) and ResNet100-ArcFace (ArcFace). We report accuracy for each convolutional feature using forward slash notation.
@@
Impact of convolutional feature maps as inputs to graph neural network. The graph neural network, $f(\mathbf{X},\mathbf{A})$ accepts a pair of inputs: features of the nodes ($\mathbf{X}$) and the adjacency matrix ($\mathbf{A}$). The adjacency matrix construction also requires features. We ablated using convolutional feature maps in conjunction with pixel intensity feature values and sensor pattern noise features. Therefore, we examined three different scenarios: (1) $\mathbf{X}\leftarrow$ pixel values, $\mathbf{A}\leftarrow$ sensor pattern noise, (2) $\mathbf{X}\leftarrow$ pixel values, $\mathbf{A}\leftarrow$ convolutional features and (3) $\mathbf{X}\leftarrow$ convolutional features, $\mathbf{A}\leftarrow$ sensor pattern noise. Pixel intensity values are simply flattened intensity values of the face images (we only used the first channel and resized them to 96$\times$96 for compactness); sensor pattern noise values are computed using the Enhanced PRNU features [40], and convolutional features are computed using two networks: (a) Inception-ResNet-V2 [59] that has 824 layers, and trained on the ImageNet dataset. We use the features from the final average pooling layer corresponding to $\mathbb{R}^{1\times 1536}$ embedding for each image. (b) ResNet100-ArcFace [60] as the face recognition model trained with ArcFace loss on the Glint360K face dataset from the InsightFace package with antelopev2 model. We used the output features of the model corresponding to $\mathbb{R}^{1\times 512}$ embedding for each image. Refer to the results obtained in Table VI.
@@
Comparison with image provenance algorithms. For this ablation study, we compared our performance with [20]. Although their objective was different than ours, their graph construction for the near-duplicates can be used as baseline against our method. Therefore, we adapted the official code provided by the original authors  [^8] https://github.com/CVRL/Scalable_Provenance/tree/master and compared with the graph building framework used in their paper. The baseline method performs the following steps. (i) GCM-based dissimilarity calculation in which they perform interest point detection on a set of near-duplicates followed by geometrically consistent matching (GCM) of the interest points. (ii) MI-based dissimilarity calculation in which they perform homography estimation from the GCM followed by mutual information (MI) calculation from the warped and color-matched patches. (iii) Construction of provenance graph in which they perform Kruskal’s Minimal Spanning Tree algorithm using the dissimilarity matrix. Refer to [20] for details. We used the SURF descriptor and similarity transformation for homography estimation. As the dissimilarity matrix from MI was asymmetric, we used the Oriented Kruskal [11] algorithm (classical Kruskal’s algorithm needs symmetric matrix). We achieved root identification and IPT reconstruction accuracy $\sim 0.1\%$. We tried different combinations of using only GCM-dissimilarity matrix, MI-dissimilarity matrix and combined (min-max normalized) dissimilarity matrix. However, we did not observe improvement in performance. Note MI captures similarity between two images, the authors in [20] use it to determine relationship between the query and the donor images that were used to generate the composite. However, in our scenario with near-duplicate faces with exactly same pose and background, implying source and target images are perfectly registered, mutual information is not discriminative enough to deduce the relationship between the near-duplicates. Our findings reaffirm that mutual information and geometric consistent matching may work on composite images originating from different donor images but fails to distinguish between subtle edits in face images.